OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + 4x - 60

OpenStudy (anonymous):

Factorise it first

OpenStudy (anonymous):

A. x - 10 B. x2 + 10 C. x + 6 D. x + 10

OpenStudy (anonymous):

Factor it...

OpenStudy (anonymous):

x^2 + 4x - 60 = x^2 - 6x + 10x - 60 = x(x - 6) + 10(x - 6) = (x - 6)(x + 10)